A metal crystallizes in bcc structure. Edge length of the unit cell is 300 pm . how many unit cells are in 5cm3 of the metal?

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Edge length = 300 pm = 3 * 10^-10 m
 
Volume of unit cell = a³ = 27 * 10^-30 m
 
Now nos. of unit cell = given volume / calculated volume = 5/27 *10^-30
 
Hope it helps